10.) If given 290mL of a diluted glucose solution 0.75% (w/v) that was made by adding 30mL of a stock glucose solution and 260mL of water, what was the original concentration of the stock glucose solution expressed as grams/100mL or % (w/v).

13.) If you have 2mL of a 2.5 X 10-7 M [H+] and it is diluted by adding 3mL of H2O, what is the new concentration of your solution?

14.) Suppose you are given a 50% (w/v) solution of sucrose and asked to prepare 250mL of a 1 to 5 dilution. How much of the stock solution should you add? How much solvent should you add? What will be the final concentration in % (w/v) of your new solution?

11.) Convert 20.9 lbs per gallon to g/mL.

12.) John’s cholesterol level is 285 mg/dL before he goes on a very stringent diet. After 6 months of dieting his cholesterol level is 170 mg/dL. What percentage drop does this represent? After 1 more month, his cholesterol levels then rise 20% (because John likes pizza). What is his most recent cholesterol level?
Answer #1

1. lets assume the stock concentration of glucose solution= C1

Volume added from the stock glucose= 30 ml

Concentration of diluted glucose solution=0.75% (w/v)

Final volume of diluted glucose solution= 290 ml

So:

C1V1=C2V2

C1 X 30=0.75 X 290

C1=7.25% (w/v)

If you have 2mL of a 2.5 X 10-7 M [H+] and it is diluted by adding 3mL of H2O, what is the new concentration of your solution?

Initial volume V1= 2mL

Initial molarity M1= 2.5 X 10-7 M

Final volume V2-=2+3= 5 ml

So

M1V1=M1V2

2.5 X 10-7 X 2=M2 X 5

M2=1 X 10-7 M

14.) Suppose you are given a 50% (w/v) solution of sucrose and asked to prepare 250mL of a 1 to 5 dilution. How much of the stock solution should you add? How much solvent should you add? What will be the final concentration in % (w/v) of your new solution?

Stock concentration= 50% (w/v)

final volume V2 = 250mL

Dilution 1 to 5

so final concentration C2=50/5= 10%(w/v)

So volume V1 required can be calculated as

C1V1=C2V2

50 X V1=10 X 250

V1= 50 ml of sucrose (50% (w/v)) and 200 ml of solvent should be added to get 10% final concentration

11.) Convert 20.9 lbs per gallon to g/mL.

Ans: 1 lbs per gallon= 0.119826 g/mL

20.9 lbs per gallon= 0.119826 X 20.9= 2.50g/mL